# Revisiting Non-Acyclic GFlowNets in Discrete Environments

### v2

@@ 1 Introduction
Generative Flow Networks (GFlowNets, Bengio et al., 2021) are models that aim to sample discrete objects from distributions known proportionally up to a constant. They operate by constructing an object through a sequence of stochastic transitions defined by a forward policy. GFlowNets have been successfully applied in various areas, starting from molecule generation (Bengio et al., 2021; Shen et al., 2024; Koziarski et al., 2024; Cretu et al., 2025) and biological sequence design (Jain et al., 2022; Kim et al., 2024) to combinatorial optimization (Zhang et al., 2023a, b; Kim et al., 2025) and fine-tuning of large language models and diffusion models (Hu et al., 2023; Venkatraman et al., 2024; Zhang et al., 2025; Uehara et al., 2024; Lee et al., 2025). The detailed theoretical foundations of GFlowNets in discrete environments were developed in (Bengio et al., 2023). While the majority of GFlowNet literature considers the discrete setting, it is possible to apply the methodology of continuous GFlowNets (Lahlou et al., 2023) to sampling problems on more general spaces.
The main idea behind the generation process in GFlowNets lies in sampling trajectories in the appropriately constructed directed acyclic graph environment instead of working directly in the object space. A standard intuition behind this process is a sequence of actions applied in order to construct a composite object from ”blocks”. One of the limitations of this setting is that it requires acyclicity. While this limitation can be naturally interpreted in, e.g., molecule generation setting, it can confine the practical design of GFlowNet environments, as well as restrict the applicability of GFlowNets in other scenarios. A motivational example for non-acyclic GFlowNets presented by (Brunswic et al., 2024) is related to modeling distributions over objects with intrinsic symmetries. Consider a class of environments where states are elements of some group, e.g. symmetric group or Rubik’s Cube group. The transitions are given via a generating set of this group, thus corresponding to applying the group operation on the current state and some element of the generating set, which leads to existence of cycles. While in some cases an acyclic environment can be designed to generate group elements, such environments of “algebraic” origin naturally contain cycles, thus falling under the area of our study. In addition, there is a growing body of work which connects GFlowNets and Reinforcement Learning (Tiapkin et al., 2024; Mohammadpour et al., 2024; Deleu et al., 2024; He et al., 2024). Most RL environments contain cycles, thus understanding how GFlowNets can be applied in cyclic environments and connecting them to RL in such case can be a crucial step towards further bridging two research fields.
@@ 2.2 GFlowNets in Non-Acyclic Environments
$\mathcal{L}_{\texttt{SDB}}(s\to s^{\prime})&\triangleq\log(1+\varepsilon\Delta^{2}(s,s^{\prime},\theta))(1+\eta\mathcal{F}_{\theta}(s))\,,\\ \ \Delta(s,s^{\prime},\theta)&\triangleq\mathcal{F}_{\theta}(s)\mathcal{P}_{\mathrm{F}}(s^{\prime}|s,\theta)-\mathcal{F}_{\theta}(s^{\prime})\mathcal{P}_{\mathrm{B}}(s|s^{\prime},\theta)\,,\end{aligned}$
@@ 3.2 Backward Policy and Trajectory Distribution
In fact, the condition $\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})>0$ together with 3.1 allows us to ensure that the sequence $s_{i}$ is a finite state-space absorbing Markov chain. Given this assumption, the proof of Lemma 3.4 almost coincides with the proof of the fact that the states of such a Markov chain are transient, see, e.g., (Douc et al., 2018). For completeness, we provide the proof in Appendix A.1.
@@ 4 Experiments
Scaling hypothesis. When $\mathcal{P}_{\mathrm{B}}$ is trainable, the main factor that plays a crucial role in loss stability in practice, i.e., controlled mean trajectory length of the trained non-acyclic GFlowNet, is the scale in which the error between flows is computed. Indeed, the standard DB loss (4) operates in log-flow scale $\Delta\log\mathcal{F}$, while standard SDB (5) operates in flow scale $\Delta\mathcal{F}$. We hypothesize that using log-flow scale losses without regularization can lead to arbitrarily large trajectory length, while flow scale losses are biased towards solutions with smaller flows and thus do not suffer from this issue.
@@ 4.1 Experimental Setting
±$\!\pm\!0.00$
DB, $\Delta\log\mathcal{F}$, $\lambda=10^{-5}$ ${0.005}$$\!\pm\!0.001$ $0.001$$\!\pm\!0.000$ $0.005$$\!\pm\!0.004$ $4.31$
±$\!\pm\!0.05$
@@ 4.2 Hypergrids
Next, we consider a larger $20\times 20\times 20\times 20$ hypergrid. An expected trajectory length induced by the chosen fixed backward policy is several orders of magnitude larger than for a smaller grid, making this approach impractical. While one can try to manually find a fixed $\mathcal{P}_{\mathrm{B}}$ with a smaller expected trajectory length, this is generally a challenging problem, thus we consider only the setting of trainable $\mathcal{P}_{\mathrm{B}}$ here. Our findings are presented in Figure 2. Similarly to $7\times 7$ grid, we find that learning in $\Delta\mathcal{F}$ scale results in a biased policy both for DB and SDB, and this bias is noticeably larger than in the smaller grid. In $\Delta\log\mathcal{F}$ scale, both DB and SDB employed with state flow regularization learn to correctly sample from the reward distribution. While all methods converge to similar expected trajectory length, $\Delta\mathcal{F}$ scale losses have smaller $n_{\tau}$ in the middle of the training even when employed without regularization, which supports our scaling hypothesis. In addition, Figure 4 in Appendix D shows that for both losses in $\Delta\log\mathcal{F}$ scale, a mean length of sampled trajectories tends to infinity when the training is done without state flow regularization. Finally, we note that $\Delta\log\mathcal{F}$ losses correctly learn the true normalizing constant $\mathcal{Z}$, while $\Delta\mathcal{F}$ losses perform worse.
@@ 4.3 Permutations
Next, we consider the environment corresponding to the Cayley Graph of the of the symmetric group $\mathrm{S}_{n}$ (group of permutations on $n$ elements $\{1,2,\dots,n\}$) from (Brunswic et al., 2024). Each state $s\in\mathcal{S}\setminus\{s_{0},s_{f}\}$ is a permutation of fixed length $(s(1),\dots,s(n))$, and there are $n-1$ possible transitions that correspond to swapping a pair of adjacent elements $s(k)$ and $s(k+1)$, plus a transition that corresponds to a circular shift of the permutation to the right $(s(n),s(1),\dots,s(n-1))$. In addition, each state has a terminating transition $s\to s_{f}$. GFlowNet reward utilized in the experiments of (Brunswic et al., 2024) is $\mathbb{I}[s(1)=1]$. We argue that this results in a fairly simple task, and a trivial forward policy exists that just applies circular shift until $1$ is in the first position. We opt for using a more complex reward distribution in our experiments and define GFlowNet reward in terms of the number of fixed points in a permutation $\mathcal{R}(s)=\exp\mathopen{}\mathclose{{}\left(\frac{1}{2}\sum_{k=1}^{n}\mathbb{I}\{s(k)=k\}}\right)$.
@@ Appendix A Proofs
Consider a random walk on $\mathcal{G}$ with reversed edges transition probabilities given by backward policy. Specifically, we define a Markov chain $\{X_{n}\}_{n=0}^{\infty}$, such that $X_{0}=s_{f}$ a.s. and $\mathbb{P}[X_{i}=s\mid X_{i-1}=s^{\prime}]=\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})$. Let also $\mathbb{P}[X_{i}=s_{0}\mid X_{i-1}=s_{0}]=1$, i.e., $s_{0}$ is an absorbing state. We want to show that
@@
    The random walk terminates at $s_{0}$ with probability 1: $\mathbb{P}[\exists n:X_{n}=s_{0}]=1$;
@@
    The expected length of a walk is finite: $\mathbb{E}[n_{\tau}]=\mathbb{E}[\sum_{n=0}^{\infty}\mathbb{I}\{X_{i}\not=s_{0}\}]<\infty$.
@@
$\sum_{\tau\in\mathcal{T}}\mathbb{P}[(X_{n_{\tau}+1},\ldots,X_{0})=\tau]=\mathbb{P}[\exists\tau\in\mathcal{T}:(X_{n_{\tau}+1},\ldots,X_{0})=\tau]=\mathbb{P}[\exists n:X_{n}=s_{0}]\,,$
@@
First consider any intermediate state $s\in\mathcal{S}\setminus\{s_{0},s_{f}\}$ and define a Markov chain $\{Y_{n}\}_{n=0}^{\infty}$ with the same transition probabilities as $\{X_{n}\}_{n=0}^{\infty}$, with $Y_{0}=s$ a.s. We define $p_{s}\triangleq\mathbb{P}[\exists n>0:Y_{n}=s]$, i.e. the probability that $\{Y_{n}\}$ returns to $s$. First, notice that $p_{s}<1$. Indeed, based on our assumptions on $\mathcal{G}$, there exists at least one path $\tau$ from $s_{0}$ to $s$, and furthermore, there exists such a path without cycles. In this case, we have
@@
by the condition on $\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})>0$ for $s\to s^{\prime}\in\mathcal{E}$. Notice that $\{(Y_{n_{\tau}+1},\ldots,Y_{0})=\tau\}\cap\{\exists n>0:Y_{n}=s\}=\emptyset$, since if the trajectory of the random walk has already reached $s_{0}$, it will never return to $s$. Thus, $p_{s}=\mathbb{P}[\exists n>0:Y_{n}=s]<1$.
Next, for each state $s\in\mathcal{S}\setminus\{s_{0},s_{f}\}$ we define $N_{s}\triangleq\sum_{i=0}^{\infty}\mathbb{I}\{X_{i}=s\}$ as the number of visits of a state $s$ encountered by the original process. Also, let us define $N^{\prime}_{s}\triangleq\sum_{i=0}^{\infty}\mathbb{I}\{Y_{i}=s\}$ as the number of visits of a state $s$ during the backward random walk that starts at $s$. We notice that
$\mathbb{E}[N_{s}]=\mathbb{P}(\exists k>0:X_{k}=s)\mathbb{E}[N_{s}^{\prime}]\leq\mathbb{E}[N_{s}^{\prime}]\,,$
@@
$\mathbb{E}[n_{\tau}]=\mathbb{E}\mathopen{}\mathclose{{}\left[\sum_{k=0}^{\infty}\mathbb{I}\{X_{k}\not=s_{0}\}}\right]=\mathbb{E}\mathopen{}\mathclose{{}\left[\sum_{s\in\mathcal{S}\setminus\{s_{0},s_{f}\}}\sum_{k=0}^{\infty}\mathbb{I}\{X_{k}=s\}}\right]=\sum_{s\in\mathcal{S}\setminus\{s_{0},s_{f}\}}\mathbb{E}[N_{s}]\leq\sum_{s\in\mathcal{S}\setminus\{s_{0},s_{f}\}}\mathbb{E}[N_{s}^{\prime}]<+\infty\,.$
@@
$\begin{split}\mathcal{F}(s\to s^{\prime})&=\mathcal{F}(s_{f})\;\mathbb{E}_{\tau}\mathopen{}\mathclose{{}\left[\sum_{t=0}^{n_{\tau}}\mathbb{I}\{s_{t}=s,s_{t+1}=s^{\prime}\}}\right]=\mathcal{F}(s_{f})\mathbb{E}_{n_{\tau}}\mathopen{}\mathclose{{}\left[\mathbb{E}_{\tau}\mathopen{}\mathclose{{}\left[\sum_{t=0}^{n_{\tau}+1}\mathbb{I}\{s_{t}=s,s_{t+1}=s^{\prime}\}\mid n_{\tau}}\right]}\right],\\ \mathcal{F}(s)&=\mathcal{F}(s_{f})\;\mathbb{E}_{\tau}\mathopen{}\mathclose{{}\left[\sum_{t=0}^{n_{\tau}+1}\mathbb{I}\{s_{t}=s\}}\right]=\mathcal{F}(s_{f})\mathbb{E}_{n_{\tau}}\mathopen{}\mathclose{{}\left[\mathbb{E}_{\tau}\mathopen{}\mathclose{{}\left[\sum_{t=0}^{n_{\tau}+1}\mathbb{I}\{s_{t}=s\}\mid n_{\tau}}\right]}\right].\end{split}$
@@
$\begin{split}\mathcal{F}(s)&=\mathcal{F}(s_{f})\mathbb{E}_{n_{\tau}}\mathopen{}\mathclose{{}\left[\mathbb{E}_{\tau}\mathopen{}\mathclose{{}\left[\sum_{t=1}^{n_{\tau}+1}\sum_{s^{\prime\prime}\in\operatorname{in}(s)}\mathbb{I}\{s_{t-1}=s^{\prime\prime},s_{t}=s\}\mid n_{\tau}}\right]}\right]\\ &=\sum_{s^{\prime\prime}\in\operatorname{in}(s)}\mathcal{F}(s_{f})\mathbb{E}_{n_{\tau}}\mathopen{}\mathclose{{}\left[\mathbb{E}_{\tau}\mathopen{}\mathclose{{}\left[\sum_{t=1}^{n_{\tau}+1}\mathbb{I}\{s_{t-1}=s^{\prime\prime},s_{t}=s\}\mid n_{\tau}}\right]}\right]=\sum_{s^{\prime\prime}\in\operatorname{in}(s)}\mathcal{F}(s^{\prime\prime}\to s).\end{split}$
@@
$\begin{split}\mathcal{F}(s)&=\mathcal{F}(s_{f})\mathbb{E}_{n_{\tau}}\mathopen{}\mathclose{{}\left[\mathbb{E}_{\tau}\mathopen{}\mathclose{{}\left[\sum_{t=0}^{n_{\tau}}\sum_{s^{\prime}\in\operatorname{out}(s)}\mathbb{I}\{s_{t}=s,s_{t+1}=s^{\prime}\}\mid n_{\tau}}\right]}\right]\\ &=\sum_{s^{\prime}\in\operatorname{out}(s)}\mathcal{F}(s_{f})\mathbb{E}_{n_{\tau}}\mathopen{}\mathclose{{}\left[\mathbb{E}_{\tau}\mathopen{}\mathclose{{}\left[\sum_{t=0}^{n_{\tau}}\mathbb{I}\{s_{t}=s,s_{t+1}=s^{\prime}\}\mid n_{\tau}}\right]}\right]=\sum_{s^{\prime}\in\operatorname{out}(s)}\mathcal{F}(s\to s^{\prime}).\\ \end{split}$
Next, we prove the key relation $\mathcal{F}(s\to s^{\prime})=\mathcal{F}(s^{\prime})\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})$. We have
$\begin{split}\mathcal{F}(s\to s^{\prime})&=\mathcal{F}(s_{f})\mathbb{E}_{n_{\tau}}\mathopen{}\mathclose{{}\left[\mathbb{E}_{\tau}\mathopen{}\mathclose{{}\left[\sum_{t=0}^{n_{\tau}}\mathbb{I}\{s_{t}=s,s_{t+1}=s^{\prime}\}\mid n_{\tau}}\right]}\right]\\ &=\mathcal{F}(s_{f})\mathbb{E}_{n_{\tau}}\mathopen{}\mathclose{{}\left[\sum_{t=0}^{n_{\tau}}\mathbb{E}_{\tau}\mathopen{}\mathclose{{}\left[\mathbb{I}\{s_{t}=s,s_{t+1}=s^{\prime}\}\mid n_{\tau}}\right]}\right]\\ &=\mathcal{F}(s_{f})\mathbb{E}_{n_{\tau}}\mathopen{}\mathclose{{}\left[\sum_{t=0}^{n_{\tau}}\mathbb{P}\mathopen{}\mathclose{{}\left(s_{t}=s,s_{t+1}=s^{\prime}\mid n_{\tau}}\right)}\right]\\ &\overset{(a)}{=}\mathcal{F}(s_{f})\mathbb{E}_{n_{\tau}}\mathopen{}\mathclose{{}\left[\sum_{t=0}^{n_{\tau}}\mathbb{P}\mathopen{}\mathclose{{}\left(s_{t+1}=s^{\prime}\mid n_{\tau}}\right)\mathbb{P}\mathopen{}\mathclose{{}\left(s_{t}=s\mid s_{t+1}=s^{\prime},n_{\tau}}\right)}\right]\\ &=\mathcal{F}(s_{f})\mathbb{E}_{n_{\tau}}\mathopen{}\mathclose{{}\left[\sum_{t=0}^{n_{\tau}}\mathbb{P}\mathopen{}\mathclose{{}\left(s_{t+1}=s^{\prime}\mid n_{\tau}}\right)}\right]\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})=\mathcal{F}(s^{\prime})\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime}).\\ \end{split}$
Here in (a) we used the Markov property of trajectory distribution. Finally, by definition, ${\mathcal{F}}(s_{0})$ is equal to, $\mathcal{F}(s_{f})$ multiplied by the expected number of times $\tau\sim\mathcal{P}$ visits $s_{0}$, where the latter is always $1$, so we have ${\mathcal{F}}(s_{0})=\mathcal{F}(s_{f})$.
@@
^$\hat{\mathcal{F}}(s\to s^{\prime})=\hat{\mathcal{F}}(s^{\prime})\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})=\frac{\hat{\mathcal{F}}(s^{\prime})}{\mathcal{F}(s^{\prime})}\mathcal{F}(s\to s^{\prime})=C(s^{\prime})\mathcal{F}(s\to s^{\prime})\,,$
@@
^$\hat{\mathcal{F}}(s)=\sum\limits_{s^{\prime}\in\operatorname{out}(s)}\hat{\mathcal{F}}(s\to s^{\prime})=\sum\limits_{s^{\prime\prime}\operatorname{in}(s)}\hat{\mathcal{F}}(s^{\prime\prime}\to s).$
@@
^$\hat{\mathcal{F}}(s)=C(s)\mathcal{F}(s)=\sum\limits_{s^{\prime}\in\operatorname{out}(s)}C(s^{\prime})\mathcal{F}(s\to s^{\prime}).$
@@
In addition, ^$\hat{\mathcal{F}}(s_{f})$ is equal to, by definition, $\mathcal{F}(s_{f})$ multiplied by the expected number of times $\tau\sim\mathcal{P}$ visits $s_{f}$, where the latter is $1$, so we have ^$\hat{\mathcal{F}}(s_{f})=\mathcal{F}(s_{f})$, thus an additional equation is $C(s_{f})=1$. In total, we have $|\mathcal{S}|$ variables and $|\mathcal{S}|$ equations, and are interested in strictly positive solutions. Firstly, there exists a trivial solution $C(s)=1$ for each $s\in\mathcal{S}$, which is an only constant solution since $C(s_{f})=1$.
@@
$11=\frac{\sum_{s^{\prime}\in\operatorname{out}(s_{t})}C(s^{\prime})\mathcal{F}(s_{t}\to s^{\prime})}{C(s_{t})\mathcal{F}(s_{t})}<\frac{\sum_{s^{\prime}\in\operatorname{out}(s_{t})}C(s_{t})\mathcal{F}(s_{t}\to s^{\prime})}{C(s_{t})\mathcal{F}(s_{t})}=\frac{\sum_{s^{\prime}\in\operatorname{out}(s_{t})}\mathcal{F}(s_{t}\to s^{\prime})}{\mathcal{F}(s_{t})}=$ 1 .

### v3

@@ 1 Introduction
Generative Flow Networks (GFlowNets, Bengio et al., 2021) are models that aim to sample discrete objects from distributions known proportionally up to a constant. They operate by constructing an object through a sequence of stochastic transitions defined by a forward policy. GFlowNets have been successfully applied in various areas, starting from molecule generation (Bengio et al., 2021; Shen et al., 2024; Koziarski et al., 2024; Cretu et al., 2025) and biological sequence design (Jain et al., 2022; Kim et al., 2024) to combinatorial optimization (Zhang et al., 2023a, b; Kim et al., 2025) and fine-tuning of large language models and diffusion models (Hu et al., 2023; Venkatraman et al., 2024; Uehara et al., 2024; Zhang et al., 2025; Lee et al., 2025). The detailed theoretical foundations of GFlowNets in discrete environments were developed in (Bengio et al., 2023). While the majority of GFlowNet literature considers the discrete setting, it is possible to apply the methodology of continuous GFlowNets (Lahlou et al., 2023) to sampling problems on more general spaces.
The main idea behind the generation process in GFlowNets lies in sampling trajectories in the appropriately constructed directed acyclic graph environment instead of working directly in the object space. A standard intuition behind this process is a sequence of actions applied in order to construct a composite object from ”blocks”. One of the limitations of this setting is that it requires acyclicity. While this limitation can be naturally interpreted in, e.g., molecule generation setting, it can confine the practical design of GFlowNet environments, as well as restrict the applicability of GFlowNets in other scenarios. A motivational example for non-acyclic GFlowNets presented by (Brunswic et al., 2024) is related to modeling distributions over objects with intrinsic symmetries. Consider a class of environments where states are elements of some group, e.g. symmetric group or Rubik’s Cube group. The transitions are given via a generating set of this group, thus corresponding to applying the group operation on the current state and some element of the generating set, which leads to the existence of cycles. While in some cases an acyclic environment can be designed to generate group elements, such environments of “algebraic” origin naturally contain cycles, thus falling under the area of our study. In addition, there is a growing body of work that connects GFlowNets and Reinforcement Learning (Tiapkin et al., 2024; Mohammadpour et al., 2024; Deleu et al., 2024; He et al., 2024). Most RL environments contain cycles, thus understanding how GFlowNets can be applied in cyclic environments and connecting them to RL in such cases can be a crucial step towards further bridging two research fields.
@@ 2.2 GFlowNets in Non-Acyclic Environments
[…]
@@ 3.2 Backward Policy and Trajectory Distribution
In fact, the condition $\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})>0$ together with Assumption 3.1 allows us to ensure that the sequence $s_{i}$ is a finite state-space absorbing Markov chain. Given this assumption, the proof of Lemma 3.4 almost coincides with the proof of the fact that the states of such a Markov chain are transient, see, e.g., (Douc et al., 2018). For completeness, we provide the proof in Appendix A.1.
@@ 4 Experiments
Scaling hypothesis. When $\mathcal{P}_{\mathrm{B}}$ is trainable, the main factor that plays a crucial role in loss stability in practice, i.e., the controlled mean trajectory length of the trained non-acyclic GFlowNet, is the scale in which the error between flows is computed. Indeed, the standard DB loss (4) operates in log-flow scale $\Delta\log\mathcal{F}$, while standard SDB (5) operates in flow scale $\Delta\mathcal{F}$. We hypothesize that using log-flow scale losses without regularization can lead to arbitrarily large trajectory length, while flow scale losses are biased towards solutions with smaller flows and thus do not suffer from this issue.
@@ 4.1 Experimental Setting
$\pm\!0.00$
DB, $\Delta\log\mathcal{F}$, $\lambda=10^{-5}$ 0.005{0.005} $\pm\!0.001$ $0.001$ $\pm\!0.000$ $0.005$ $\pm\!0.004$ $4.31$
$\pm\!0.05$
@@ 4.2 Hypergrids
Next, we consider a larger $20\times 20\times 20\times 20$ hypergrid. An expected trajectory length induced by the chosen fixed backward policy is several orders of magnitude larger than for a smaller grid, making this approach impractical. While one can try to manually find a fixed $\mathcal{P}_{\mathrm{B}}$ with a smaller expected trajectory length, this is generally a challenging problem, thus, we consider only the setting of trainable $\mathcal{P}_{\mathrm{B}}$ here. Our findings are presented in Figure 2. Similarly to $7\times 7$ grid, we find that learning in $\Delta\mathcal{F}$ scale results in a biased policy both for DB and SDB, and this bias is noticeably larger than in the smaller grid. In $\Delta\log\mathcal{F}$ scale, both DB and SDB employed with state flow regularization learn to correctly sample from the reward distribution. While all methods converge to similar expected trajectory length, $\Delta\mathcal{F}$ scale losses have smaller $n_{\tau}$ in the middle of the training even when employed without regularization, which supports our scaling hypothesis. In addition, Figure 4 in Appendix D shows that for both losses in $\Delta\log\mathcal{F}$ scale, a mean length of sampled trajectories tends to infinity when the training is done without state flow regularization. Finally, we note that $\Delta\log\mathcal{F}$ losses correctly learn the true normalizing constant $\mathcal{Z}$, while $\Delta\mathcal{F}$ losses perform worse.
@@ 4.3 Permutations
Next, we consider the environment corresponding to the Cayley Graph of the symmetric group $\mathrm{S}_{n}$ (group of permutations on $n$ elements $\{1,2,\dots,n\}$) from (Brunswic et al., 2024). Each state $s\in\mathcal{S}\setminus\{s_{0},s_{f}\}$ is a permutation of fixed length $(s(1),\dots,s(n))$, and there are n1n-1 possible transitions that correspond to swapping a pair of adjacent elements $s(k)$ and $s(k+1)$, plus a transition that corresponds to a circular shift of the permutation to the right $(s(n),s(1),\dots,s(n-1))$. In addition, each state has a terminating transition $s\to s_{f}$. GFlowNet reward utilized in the experiments of (Brunswic et al., 2024) is $\mathbb{I}[s(1)=1]$. We argue that this results in a fairly simple task, and a trivial forward policy exists that just applies circular shift until $1$ is in the first position. We opt for using a more complex reward distribution in our experiments and define GFlowNet reward in terms of the number of fixed points in a permutation $\mathcal{R}(s)=\exp\mathopen{}\mathclose{{\left(\frac{1}{2}\sum_{k=1}^{n}\mathbb{I}\{s(k)=k\}}}\right)$.
@@ Appendix A Proofs
Consider a random walk on $\mathcal{G}$ with reversed edges transition probabilities given by backward policy. Specifically, we define a Markov chain $\{X_{n}\}_{n=0}^{\infty}$, such that $X_{0}=s_{f}$ a.s. and $\mathbb{P}[X_{t}=s\mid X_{t-1}=s^{\prime}]=\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})$. Let also $\mathbb{P}[X_{t}=s_{0}\mid X_{t-1}=s_{0}]=1$, i.e., $s_{0}$ is an absorbing state. We want to show that
@@
    The random walk terminates at $s_{0}$ with probability 1: $\mathbb{P}[\exists t:X_{t}=s_{0}]=1$;
@@
    The expected length of a walk is finite: $\mathbb{E}[n_{\tau}]=\mathbb{E}[\sum_{t=0}^{\infty}\mathbb{I}\{X_{t}\not=s_{0}\}]<\infty$.
@@
$\sum_{\tau\in\mathcal{T}}\mathbb{P}[(X_{n_{\tau}+1},\ldots,X_{0})=\tau]=\mathbb{P}[\exists\tau\in\mathcal{T}:(X_{n_{\tau}+1},\ldots,X_{0})=\tau]=\mathbb{P}[\exists t:X_{t}=s_{0}]\,$,
@@
First consider any intermediate state $s\in\mathcal{S}\setminus\{s_{0},s_{f}\}$ and define a Markov chain $\{Y_{n}\}_{n=0}^{\infty}$ with the same transition probabilities as $\{X_{n}\}_{n=0}^{\infty}$, with $Y_{0}=s$ a.s. We define $p_{s}\triangleq\mathbb{P}[\exists t>0:Y_{t}=s]$, i.e. the probability that $\{Y_{n}\}$ returns to $s$. First, notice that $p_{s}<1$. Indeed, based on our assumptions on $\mathcal{G}$, there exists at least one path $\tau$ from $s_{0}$ to $s$, and furthermore, there exists such a path without cycles. In this case, we have
@@
by the condition on $\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})>0$ for $s\to s^{\prime}\in\mathcal{E}$. Notice that $\{(Y_{n_{\tau}+1},\ldots,Y_{0})=\tau\}\cap\{\exists t>0:Y_{t}=s\}=\emptyset$, since if the trajectory of the random walk has already reached $s_{0}$, it will never return to $s$. Thus, $p_{s}=\mathbb{P}[\exists t>0:Y_{t}=s]<1$.
Next, for each state $s\in\mathcal{S}\setminus\{s_{0},s_{f}\}$ we define $N_{s}\triangleq\sum_{t=0}^{\infty}\mathbb{I}\{X_{t}=s\}$ as the number of visits of a state $s$ encountered by the original process. Also, denote by $N^{\prime}_{s}\triangleq\sum_{t=0}^{\infty}\mathbb{I}\{Y_{t}=s\}$ the number of visits of a state $s$ during the backward random walk that starts at $s$. We notice that
$\mathbb{E}[N_{s}]=\mathbb{P}(\exists t>0:X_{t}=s)\mathbb{E}[N_{s}^{\prime}]\leq\mathbb{E}[N_{s}^{\prime}]\,$,
@@
$\mathbb{E}[n_{\tau}]=\mathbb{E}\mathopen{}\mathclose{{\left[\sum_{t=0}^{\infty}\mathbb{I}\{X_{t}\not=s_{0}\}}}\right]=\mathbb{E}\mathopen{}\mathclose{{\left[\sum_{s\in\mathcal{S}\setminus\{s_{0},s_{f}\}}\sum_{t=0}^{\infty}\mathbb{I}\{X_{t}=s\}}}\right]=\sum_{s\in\mathcal{S}\setminus\{s_{0},s_{f}\}}\mathbb{E}[N_{s}]\leq\sum_{s\in\mathcal{S}\setminus\{s_{0},s_{f}\}}\mathbb{E}[N_{s}^{\prime}]<+\infty\,.$
@@
$\mathcal{F}(s\to s^{\prime})=\mathcal{F}(s_{f})\;\mathbb{E}_{\tau}\mathopen{}\mathclose{{\left[\sum_{t=0}^{n_{\tau}}\mathbb{I}\{s_{t}=s,s_{t+1}=s^{\prime}\}}}\right],\quad\mathcal{F}(s)=\mathcal{F}(s_{f})\;\mathbb{E}_{\tau}\mathopen{}\mathclose{{\left[\sum_{t=0}^{n_{\tau}+1}\mathbb{I}\{s_{t}=s\}}}\right].$
@@
$\begin{split}\mathcal{F}(s)&=\mathcal{F}(s_{f})\mathbb{E}_{\tau}\mathopen{}\mathclose{{\left[\sum_{t=1}^{n_{\tau}+1}\sum_{s^{\prime\prime}\in\operatorname{in}(s)}\mathbb{I}\{s_{t-1}=s^{\prime\prime},s_{t}=s\}}}\right]\\ &=\sum_{s^{\prime\prime}\in\operatorname{in}(s)}\mathcal{F}(s_{f})\mathbb{E}_{\tau}\mathopen{}\mathclose{{\left[\sum_{t=1}^{n_{\tau}+1}\mathbb{I}\{s_{t-1}=s^{\prime\prime},s_{t}=s\}}}\right]=\sum_{s^{\prime\prime}\in\operatorname{in}(s)}\mathcal{F}(s^{\prime\prime}\to s).\end{split}$
@@
[…]
Next, we prove the key relation $\mathcal{F}(s\to s^{\prime})=\mathcal{F}(s^{\prime})\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})$. Let $\{X_{n}\}_{n=0}^{\infty}$ be the Markov chain defined in the proof of Lemma 3.4, corresponding to a random walk on $\mathcal{G}$ with reversed edges and transition probabilities given by $\mathcal{P}_{\mathrm{B}}$. We have
[…]
Here in (a) we used the fact that
$\mathbb{E}\mathopen{}\mathclose{{\left[\sum_{t=0}^{+\infty}\mathbb{I}\{X_{t}=s^{\prime},X_{t+1}=s\}}}\right]\leq\mathbb{E}[n_{\tau}]<\infty$
due to Lemma 3.4, so the expectation is finite and we can apply Fubini’s theorem. In (b) we used the Markov property of $\{X_{n}\}_{n=0}^{\infty}$. Finally, ${\mathcal{F}}(s_{0})=\mathcal{F}(s_{f})$ by definition (8), since each trajectory $\tau\sim\mathcal{P}$ visits $s_{0}$ exactly once.
@@
$\hat{\mathcal{F}}(s\to s^{\prime})=\hat{\mathcal{F}}(s^{\prime})\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})=\frac{\hat{\mathcal{F}}(s^{\prime})}{\mathcal{F}(s^{\prime})}\mathcal{F}(s\to s^{\prime})=C(s^{\prime})\mathcal{F}(s\to s^{\prime})\,$,
@@
$\hat{\mathcal{F}}(s)=\sum\limits_{s^{\prime}\in\operatorname{out}(s)}\hat{\mathcal{F}}(s\to s^{\prime})=\sum\limits_{s^{\prime\prime}\operatorname{in}(s)}\hat{\mathcal{F}}(s^{\prime\prime}\to s).$
@@
$\hat{\mathcal{F}}(s)=C(s)\mathcal{F}(s)=\sum\limits_{s^{\prime}\in\operatorname{out}(s)}C(s^{\prime})\mathcal{F}(s\to s^{\prime}).$
@@
In addition, $\hat{\mathcal{F}}(s_{f})$ is equal to, by definition, $\mathcal{F}(s_{f})$ multiplied by the expected number of times $\tau\sim\mathcal{P}$ visits $s_{f}$, where the latter is $1$, so we have $\hat{\mathcal{F}}(s_{f})=\mathcal{F}(s_{f})$, thus an additional equation is $C(s_{f})=1$. In total, we have $|\mathcal{S}|$ variables and $|\mathcal{S}|$ equations, and are interested in strictly positive solutions. Firstly, there exists a trivial solution $C(s)=1$ for each $s\in\mathcal{S}$, which is an only constant solution since $C(s_{f})=1$.
@@
$1=\frac{\sum_{s^{\prime}\in\operatorname{out}(s_{t})}C(s^{\prime})\mathcal{F}(s_{t}\to s^{\prime})}{C(s_{t})\mathcal{F}(s_{t})}<\frac{\sum_{s^{\prime}\in\operatorname{out}(s_{t})}C(s_{t})\mathcal{F}(s_{t}\to s^{\prime})}{C(s_{t})\mathcal{F}(s_{t})}=\frac{\sum_{s^{\prime}\in\operatorname{out}(s_{t})}\mathcal{F}(s_{t}\to s^{\prime})}{\mathcal{F}(s_{t})}=1.$